The common stock of Leaning Tower of Pita, Inc., a restaurant chain, will generate the following payoffs to investors next year:
Probability Dividend Stock Price
Boom 0.35 $7 $200
Normal economy 0.45 $5 $127
Recession 0.2 $1 $7
The company goes out of business if a recession hits. Calculate the expected rate of return and standard deviation of return to Leaning Tower of Pita share-holders. The stock is selling today for $95. (Round your answers to 2 decimal places.)
Expected rate of return         %
Standard deviation      %
Answer #1

Returns:
Boom=(200+7)/95-1=117.895%
Normal=(127+5)/95-1=38.947%
Recession=(7+1)/95-1=-91.579%

Expected rate of return=0.35*117.895%+0.45*38.947%+0.2*(-91.579%)=40.474%

Standard Deviation=sqrt(0.35*(117.895%-40.474%)^2+0.45*(38.947%-40.474%)^2+0.2*(-91.579%-40.474%)^2)=74.743%